In 2017-18, the Pakistan Bureau of Statistics announced that of all adult Pakistanis, 59,786,200 were employed, 3,613,800 were unemployed, and 58,800,000 were not in the labor force. Use this information to calculate the unemployment rate



1
Expert's answer
2021-05-07T09:54:56-0400

The labor force in the economy:

Labor force  =employed + unemployed

"=59,786,200+3,613,800=63,400,000"

Unemployment rate:

"=\frac{Number of unemployed}{labor force}×100"

"\frac{3,613,800}{63,400,000}×100"


"5.7\%"